OpenStudy (anonymous):

The sum of two consecutive even integers is 226. What is the larger integer?

OpenStudy (anonymous):

114

OpenStudy (anonymous):

How do you find that out?

OpenStudy (jamesj):

@mainaknag and others. I strongly encourage you not to give answers like that without explaining. Even better, show the questioner the first steps towards the answer.

OpenStudy (anonymous):

lets call the 1st integer n, so the next integer will be n+1, so n+(n+2)=226 so 2n+2=226, 2n=224, so n=112, but we want the larger one, n+2=114

OpenStudy (anonymous):

i meant to say lets call the 1st even integer n, and the next even integer n+2*